# Statistical physics analysis of graph neural networks: Approaching optimality in the contextual stochastic block model

### v2

@@ I.2.2 Analyzable model of synthetic data:
A key aspect in statistical physics is the thermodynamic limit, how should $N$ and $M$ scale together. In statistical physics we always aim at a scaling in which quantities of interest concentrate around deterministic values, and the performance of the system ranges between as bad as random guessing to as good as perfect learning. As we will see, these two requirements are satisfied in the high-dimensional limit $N\to\infty$ and $M\to\infty$ with $\alpha=N/M$ of order one. This scaling limit also aligns well with the common graph datasets that are of interest in practice, for instance Cora [30] ($N=3.10^{3}$ and $M=3.10^{3}$), DBLP [31] ($N=2.10^{4}$ and $M=2.10^{3}$), CiteSeer [32] ($N=4.10^{3}$ and $M=3.10^{3}$) and PubMed [33] ($N=2.10^{4}$ and $M=5.10^{2}$).
@@ II.2 Analyzed architecture
where $\hat{R}$ stands either for the train set $R$ or the test set $R^{\prime}$ and the expectation is taken over $y$, $u$, $A$, $X$, $R$ and $R^{\prime}$. $\mathrm{Acc}_{\mathrm{train/test}}$ is the proportion of train/test nodes that are correctly classified. A main part of the present work is dedicated to the derivation of exact expressions for the errors and the accuracies. We will then search for the architecture of the GCN that maximizes the test accuracy $\mathrm{Acc}_{\mathrm{test}}$.
@@
$c_{k}$, $c$, $t$ residual connection strengths, diffusion time
@@ III.1.1 Asymptotic characterization
$\qquad{}+\log\mathcal{N}\left(h_{0}\left|\sqrt{\mu}ym_{w}+\sqrt{Q_{w}}\zeta;V_{w}\right.\right)+\log\mathcal{N}\left(h_{>0}\left|c\odot h_{<K}+\lambda ym+Q^{1/2}\chi;V\right.\right)\ ,$
@@
[…]
@@ III.1.2 Analytical solution
$\qquad\left.{}+\sum_{l=0}^{k}\left(1+\rho\sum_{l^{\prime}=1}^{K-1-l}\lambda^{2l^{\prime}}+\frac{\alpha^{2}r^{2}}{\rho}m_{l}^{2}\right)\right)\,.$
@@ III.2.1 Asymptotic characterization
are vectors of size $K+1$, where $y=\pm 1$ is Rademacher and $\zeta_{q}\sim\mathcal{N}(0,I_{K+1})$, $\zeta_{h}\sim\mathcal{N}(0,I_{K+1})$ and $\chi\sim\mathcal{N}(0,1)$ are standard Gaussians. $\bar{s}$ determines whether the loss is active $\bar{s}=1$ or not $\bar{s}=0$. We assumed that $\ell$ is quadratic. Later we will take the limit $r\to\infty$ where $h$ is small and where $\ell$ can effectively be expanded around 00 as a quadratic potential. Notice that in the case $\delta_{\mathrm{e}}=0$ we recover the potential $\psi_{h}$ eq. (30) of the previous part.
@@
($\ -\mathrm{i}V_{qh}&\cdot\end{smallmatrix}\right)=t\mathcal{P}\left(G_{0}^{-1}D_{qh}G^{-1}\right)$ (82)
@@ III.2.2 Expansion around large regularization $r$ and continuous limit
In the limit of large regularization $h$ and $w$ are of order $1/r$; the parameters $m_{w}$, $m$, $V_{w}$ and $V$ are of order $1/r$ and $Q_{w}$ and $Q$ are of order $1/r^{2}$, while all their conjugates, $Q_{qh}$ and $V_{qh}$ are of order one. Consequently we have $G_{0}^{-1}\sim r\gg G_{h}\sim 1$ and we expand $G^{-1}$ around $G_{0}$:
@@
($\ -\mathrm{i}V_{qh}&\cdot\end{smallmatrix}\right)$ $=tD_{qh}^{-1,T}\ ,$ (87)
@@
$\left(\begin{smallmatrix}\hat{m}_{w}\\ \hat{m}\end{smallmatrix}\right)$ $=\left(\begin{smallmatrix}K\sqrt{\mu}&0\\ 0&\lambda^{\mathrm{e}}tI_{K}\cdot\end{smallmatrix}\right)D_{qh}^{-1,T}\left(\begin{smallmatrix}\hat{m}\\ \rho\end{smallmatrix}\right)$ (90)
@@
($\ -\mathrm{i}Q_{qh}&\cdot\end{smallmatrix}\right)$ $=tD_{qh}^{-1,T}\left[\left(t\delta_{\mathrm{e}}\left(\begin{smallmatrix}0&-\mathrm{i}Q_{qh}\\ 0&0\end{smallmatrix}\right)+\left(\begin{smallmatrix}\hat{m}\\ \rho\end{smallmatrix}\right)\left(\begin{smallmatrix}K\sqrt{\mu}m_{w}\\ \lambda^{\mathrm{e}}tm\end{smallmatrix}\right)^{T}\right)D_{qh}^{-1,T}+\left(\left(\begin{smallmatrix}\hat{Q}_{h}&0\\ 0&0\end{smallmatrix}\right)+\rho\left(\begin{smallmatrix}\hat{m}\\ 1\end{smallmatrix}\right)^{\otimes 2}+(1-\rho)\left(\begin{smallmatrix}\hat{m}\\ 0\end{smallmatrix}\right)^{\otimes 2}\right)D_{qh}^{-1}G_{0}D_{qh}^{-1,T}\right]$ (93)
@@
where we multiplied both sides by $D_{qh}^{T}$ and took $V_{qh}(x,z)$ for $-\mathrm{i}V_{qh}$. The solution to this integro-differential equation is
@@
$D_{qh}^{-1}(x,z)$ $=D_{qh}^{-1,T}(z,x)=\left\{\begin{array}[]{cc}\theta(x-z)&\mathrm{if}\,\delta_{\mathrm{e}}=0\\ \frac{1}{t}V_{qh}(z,x)&\mathrm{if}\,\delta_{\mathrm{e}}=1\end{array}\right.\ .$ (101)
@@
$\varphi(x)$ =$=\left\{\begin{array}[]{cc}e^{\lambda^{\mathrm{e}}tx}&\mathrm{if}\,\delta_{\mathrm{e}}=0\\ \sum_{\nu>0}^{\infty}\nu(\lambda^{\mathrm{e}})^{\nu-1}\frac{I_{\nu}(2tx)}{tx}&\mathrm{if}\,\delta_{\mathrm{e}}=1\end{array}\right.\ ,$ (106)
$\Phi(x,z)$ =$=\left\{\begin{array}[]{cc}I_{0}(2t\sqrt{xz})&\mathrm{if}\,\delta_{\mathrm{e}}=0\\ \frac{I_{1}(2t(x+z))}{t(x+z)}&\mathrm{if}\,\delta_{\mathrm{e}}=1\end{array}\right.\ .$ (109)
@@
$\quad{}+\frac{t}{\lambda^{\mathrm{e}}}\int_{0}^{x}\mathrm{d}x^{\prime}\int_{0}^{1}\mathrm{d}x^{\prime\prime}\,\varphi(x-x^{\prime})V_{h}(x^{\prime},x^{\prime\prime})\hat{m}(x^{\prime\prime})$
@@
$\qquad\left.{}+\int_{0,0^{-}}^{1^{+},1}\mathrm{d}x^{\prime\prime}\mathrm{d}z^{\prime\prime}\,\left(\hat{Q}_{h}(1-x^{\prime},x^{\prime\prime})+\mathcal{P}(\hat{m}^{\otimes 2})(1-x^{\prime},x^{\prime\prime})\right)D_{qh}^{-1}(x^{\prime\prime},z^{\prime\prime})G_{0}(z^{\prime\prime},z^{\prime})\right]$
@@
$\qquad{}+\int_{0^{-},0}^{1,1^{+}}\mathrm{d}x^{\prime\prime}\mathrm{d}x^{\prime\prime\prime}\,G_{0}(x^{\prime},x^{\prime\prime})D_{qh}^{-1,T}(x^{\prime\prime},x^{\prime\prime\prime})\left(t\delta_{\mathrm{e}}Q_{qh}(x^{\prime\prime\prime},z^{\prime})+\mathcal{P}(\hat{m})(x^{\prime\prime\prime})(\lambda^{\mathrm{e}}tm(z^{\prime})+\sqrt{\mu}m_{w}\delta(z^{\prime}))\right)$
$\qquad{}+\int_{0,0^{-}}^{1^{+},1}\mathrm{d}z^{\prime\prime}\mathrm{d}z^{\prime\prime\prime}\,\left(t\delta_{\mathrm{e}}Q_{qh}(z^{\prime\prime\prime},x^{\prime})+(\lambda^{\mathrm{e}}tm(x^{\prime})+\sqrt{\mu}m_{w}\delta(x^{\prime}))\mathcal{P}(\hat{m})(z^{\prime\prime\prime})\right)D_{qh}^{-1}(z^{\prime\prime\prime},z^{\prime\prime})G_{0}(z^{\prime\prime},z^{\prime})$
$\quad\left.{}+\int_{0^{-},0,0,0^{-}}^{1,1^{+},1^{+},1}\mathrm{d}x^{\prime\prime}\mathrm{d}x^{\prime\prime\prime}\mathrm{d}z^{\prime\prime}\mathrm{d}z^{\prime\prime\prime}\,G_{0}(x^{\prime},x^{\prime\prime})D_{qh}^{-1,T}(x^{\prime\prime},x^{\prime\prime\prime})\left(\hat{Q}_{h}(x^{\prime\prime\prime},z^{\prime\prime\prime})+\mathcal{P}(\hat{m}^{\otimes 2})(x^{\prime\prime\prime},z^{\prime\prime\prime})\right)D_{qh}^{-1}(z^{\prime\prime\prime},z^{\prime\prime})G_{0}(z^{\prime\prime},z^{\prime})\right]\ ;$
@@
and take $Q_{qh}(x,z)$ for $-\mathrm{i}Q_{qh}$. The accuracies are, with $\bar{s}=1$ for train and $\bar{s}=0$ for test:
@@
$\quad\frac{1}{2}\left(1+\mathrm{erf}\left(\frac{m(1)+(\bar{s}-\rho)V_{h}(1,1)}{\sqrt{2}\sqrt{Q_{h}(1,1)-m(1)^{2}-\rho(1-\rho)V_{h}(1,1)^{2}}}\right)\right)\ .$
@@ III.2.3 Consequences
These two limits show that at finite time $t$ the GCN avoids oversmoothing and interpolates between an estimator that is only function of the features at $t=0$ and an estimator only function of the graph at $t=\infty$. $t$ has to be fine-tuned to reach the best trade-off $t^{*}$ and the optimal performance.
@@ Acknowledgments
We acknowledge useful discussions with J. Zavatone-Veth, F. Zamponi and V. Erba. This work is supported by the Swiss National Science Foundation under grant SNFS SMArtNet (grant number 212049).
@@ Appendix A Asymptotic characterisation of the discrete GCN
The free energy of the problem is $-\beta Nf=\partial_{n}\mathbb{E}_{u,\Xi,W,y}Z^{n}(n=0)$ where the partition function is
@@ A.1 Derivation of the free energy
$\quad\quad\prod_{i}\mathcal{N}\left(h_{i,>0}\left|c\odot h_{i,<K}+y_{i}\frac{\lambda}{N}\sum_{j}y_{j}h_{j,<K};\tilde{Q}\right.\right)\prod_{i}\mathcal{N}\left(h_{i,0}\left|y_{i}\frac{\sqrt{\mu}}{N}\sum_{\nu}u_{\nu}w_{\nu};\frac{1}{N}\sum_{\nu}w_{\nu}w_{\nu}^{T}\right.\right)\ .$
@@
$\qquad\qquad{}+\log\mathcal{N}\left(h_{>0}\left|c\odot h_{<K}+\lambda ym_{<K};\tilde{Q}\right.\right)+\log\mathcal{N}\left(h_{0}\left|\sqrt{\mu}ym_{w};Q_{w}\right.\right)\ .$ (145)
@@
$I$ is the $n\times n$ identity and $J$ is the $n\times n$ matrix filled with ones. We introduce the $K\times K$ symmetric matrices $Q$ and $V$, filled with $(Q_{k})_{0\leq k\leq K-1}$ and $(V_{k})_{0\leq k\leq K-1}$ on the diagonal, and $(Q_{k,l})_{0\leq k<l\leq K-1}$ and $(V_{k,l})_{0\leq k<l\leq K-1}$ off the diagonal, such that $\tilde{Q}$ can be written in terms of Kronecker products as
@@
$\quad\quad{}+\log\mathcal{N}\left(h_{0}\left|\sqrt{\mu}ym_{w}+\sqrt{Q_{w}}\zeta;V_{w}\right.\right)+\log\mathcal{N}\left(h_{>0}\left|c\odot h_{<K}+\lambda ym+Q^{1/2}\chi;V\right.\right)\ ;$
@@
::=:= $\ \int\mathrm{d}\Theta\mathrm{d}\hat{\Theta}e^{N\phi^{(n)}(\Theta,\hat{\Theta})}\ .$ (160)
@@
$=$ $\operatorname*{extr}_{\Theta,\hat{\Theta}}\frac{\partial}{\partial n}(n=0)\phi^{(n)}(\Theta,\hat{\Theta})$ (162)
::=:= $\operatorname*{extr}_{\Theta,\hat{\Theta}}\phi(\Theta,\hat{\Theta})\ ,$ (163)
@@
$\hat{m}\to\beta\hat{m}\ ,$ $\hat{Q}\to\beta^{2}\hat{Q}\ ,$ $\hat{V}\to\beta\hat{V}\ ,$ $V\to\beta^{-1}V$ (167)
@@
[…]

### v3

@@ I.2.2 Analyzable model of synthetic data:
A key aspect in statistical physics is the thermodynamic limit, how should $N$ and $M$ scale together. In statistical physics we always aim at a scaling in which quantities of interest concentrate around deterministic values, and the performance of the system ranges between as bad as random guessing to as good as perfect learning. As we will see, these two requirements are satisfied in the high-dimensional limit $N\to\infty$ and $M\to\infty$ with $\alpha=N/M$ of order one. This scaling limit also aligns well with the common graph datasets that are of interest in practice, for instance Cora [30] ($N=3.10^{3}$ and $M=3.10^{3}$), Coauthor CS [31] ($N=2.10^{4}$ and $M=7.10^{3}$), CiteSeer [32] ($N=4.10^{3}$ and $M=3.10^{3}$) and PubMed [33] ($N=2.10^{4}$ and $M=5.10^{2}$).
@@ II.2 Analyzed architecture
where $\hat{R}$ stands either for the train set $R$ or the test set $R^{\prime}$ and the expectation is taken over $y$, $u$, AA, $X$, $R$ and $R^{\prime}$. $\mathrm{Acc}_{\mathrm{train/test}}$ is the proportion of train/test nodes that are correctly classified. A main part of the present work is dedicated to the derivation of exact expressions for the errors and the accuracies. We will then search for the architecture of the GCN that maximizes the test accuracy $\mathrm{Acc}_{\mathrm{test}}$.
@@
$_{k}, cc$, $t$ residual connection strengths, diffusion time
@@ III.1.1 Asymptotic characterization
[…]
@@
$\nabla\nabla$ being the Hessian with respect to $h$. Last, for compactness we introduce the operator $\mathcal{P}$ that, for a function $g$ in $h$, acts according to
@@ III.1.2 Analytical solution
$\hskip 18.49988pt\left.{}+\sum_{l=0}^{k}\left(1+\rho\sum_{l^{\prime}=1}^{K-1-l}\lambda^{2l^{\prime}}+\frac{\alpha^{2}r^{2}}{\rho}m_{l}^{2}\right)\right)\,.$
@@ III.2.1 Asymptotic characterization
are vectors of size $K+1$, where $y=\pm 1$ is Rademacher and $\zeta_{q}\sim\mathcal{N}(0,I_{K+1})$, $\zeta_{h}\sim\mathcal{N}(0,I_{K+1})$ and $\chi\sim\mathcal{N}(0,1)$ are standard Gaussians. $\bar{s}$ determines whether the loss is active $\bar{s}=1$ or not $\bar{s}=0$. We assumed that $\ell$ is quadratic. Later we will take the limit $r\to\infty$ where $h$ is small and where $\ell$ can effectively be expanded around 0 as a quadratic potential. Notice that in the case $\delta_{\mathrm{e}}=0$ we recover the potential $\psi_{h}$ eq. (30) of the previous part.
@@
$\left(\begin{smallmatrix}\cdot&\cdot\\ -\mathrm{i}V_{qh}&\cdot\end{smallmatrix}\right)=t\mathcal{P}\left(G_{0}^{-1}D_{qh}G^{-1}\right)$ (82)
@@ III.2.2 Expansion around large regularization $r$ and continuous limit
In the limit of large regularization $h$ and $w$ are of order $1/r$; the parameters $_{w}, mm$, $V_{w}$ and $V$ are of order $1/r$ and $Q_{w}$ and $Q$ are of order $1/r^{2}$, while all their conjugates, $Q_{qh}$ and $V_{qh}$ are of order one. Consequently we have $G_{0}^{-1}\sim r\gg G_{h}\sim 1$ and we expand $G^{-1}$ around $G_{0}$:
@@
$\left(\begin{smallmatrix}\cdot&\cdot\\ -\mathrm{i}V_{qh}&\cdot\end{smallmatrix}\right)$ $=tD_{qh}^{-1,T}\ ,$ (87)
@@
[…]
@@
$\left(\begin{smallmatrix}\cdot&\cdot\\ -\mathrm{i}Q_{qh}&\cdot\end{smallmatrix}\right)$ $=tD_{qh}^{-1,T}\left[\left(t\delta_{\mathrm{e}}\left(\begin{smallmatrix}0&-\mathrm{i}Q_{qh}\\ 0&0\end{smallmatrix}\right)+\left(\begin{smallmatrix}\hat{m}\\ \rho\end{smallmatrix}\right)\left(\begin{smallmatrix}K\sqrt{\mu}m_{w}\\ \lambda^{\mathrm{e}}tm\end{smallmatrix}\right)^{T}\right)D_{qh}^{-1,T}+\left(\left(\begin{smallmatrix}\hat{Q}_{h}&0\\ 0&0\end{smallmatrix}\right)+\rho\left(\begin{smallmatrix}\hat{m}\\ 1\end{smallmatrix}\right)^{\otimes 2}+(1-\rho)\left(\begin{smallmatrix}\hat{m}\\ 0\end{smallmatrix}\right)^{\otimes 2}\right)D_{qh}^{-1}G_{0}D_{qh}^{-1,T}\right]$ (93)
@@
where we multiplied both sides by $D_{qh}^{T}$ and took $V_{qh}(x,z)$ for $\mathrm{i}V_{qh}$. The solution to this integro-differential equation is
@@
$D_{qh}^{-1}(x,z)$ $=D_{qh}^{-1,T}(z,x)=\left\{\begin{array}[]{cc}\theta(x-z)&\mathrm{if}\,\delta_{\mathrm{e}}=0\\ \frac{1}{t}V_{qh}(z,x)&\mathrm{if}\,\delta_{\mathrm{e}}=1\end{array}\right.\$. (101)
@@
$\varphi(x)$ $=\left\{\begin{array}[]{cc}e^{\lambda^{\mathrm{e}}tx}&\mathrm{if}\,\delta_{\mathrm{e}}=0\\ \sum_{\nu>0}^{\infty}\nu(\lambda^{\mathrm{e}})^{\nu-1}\frac{I_{\nu}(2tx)}{tx}&\mathrm{if}\,\delta_{\mathrm{e}}=1\end{array}\right$.\ , (106)
$\Phi(x,z)$ $=\left\{\begin{array}[]{cc}I_{0}(2t\sqrt{xz})&\mathrm{if}\,\delta_{\mathrm{e}}=0\\ \frac{I_{1}(2t(x+z))}{t(x+z)}&\mathrm{if}\,\delta_{\mathrm{e}}=1\end{array}\right.\$. (109)
@@
$\qquad{}+\frac{t}{\lambda^{\mathrm{e}}}\int_{0}^{x}\mathrm{d}x^{\prime}\int_{0}^{1}\mathrm{d}x^{\prime\prime}\,\varphi(x-x^{\prime})V_{h}(x^{\prime},x^{\prime\prime})\hat{m}(x^{\prime\prime})$
@@
$\hskip 18.49988pt\left.{}+\int_{0,0^{-}}^{1^{+},1}\mathrm{d}x^{\prime\prime}\mathrm{d}z^{\prime\prime}\,\left(\hat{Q}_{h}(1-x^{\prime},x^{\prime\prime})+\mathcal{P}(\hat{m}^{\otimes 2})(1-x^{\prime},x^{\prime\prime})\right)D_{qh}^{-1}(x^{\prime\prime},z^{\prime\prime})G_{0}(z^{\prime\prime},z^{\prime})\right]$
@@
$\hskip 18.49988pt{}+\int_{0^{-},0}^{1,1^{+}}\mathrm{d}x^{\prime\prime}\mathrm{d}x^{\prime\prime\prime}\,G_{0}(x^{\prime},x^{\prime\prime})D_{qh}^{-1,T}(x^{\prime\prime},x^{\prime\prime\prime})\left(t\delta_{\mathrm{e}}Q_{qh}(x^{\prime\prime\prime},z^{\prime})+\mathcal{P}(\hat{m})(x^{\prime\prime\prime})(\lambda^{\mathrm{e}}tm(z^{\prime})+\sqrt{\mu}m_{w}\delta(z^{\prime}))\right)$
$\hskip 18.49988pt{}+\int_{0,0^{-}}^{1^{+},1}\mathrm{d}z^{\prime\prime\prime}\mathrm{d}z^{\prime\prime}\,\left(t\delta_{\mathrm{e}}Q_{qh}(z^{\prime\prime\prime},x^{\prime})+(\lambda^{\mathrm{e}}tm(x^{\prime})+\sqrt{\mu}m_{w}\delta(x^{\prime}))\mathcal{P}(\hat{m})(z^{\prime\prime\prime})\right)D_{qh}^{-1}(z^{\prime\prime\prime},z^{\prime\prime})G_{0}(z^{\prime\prime},z^{\prime})$
$\qquad\left.{}+\int_{0^{-},0,0,0^{-}}^{1,1^{+},1^{+},1}\mathrm{d}x^{\prime\prime}\mathrm{d}x^{\prime\prime\prime}\mathrm{d}z^{\prime\prime\prime}\mathrm{d}z^{\prime\prime}\,G_{0}(x^{\prime},x^{\prime\prime})D_{qh}^{-1,T}(x^{\prime\prime},x^{\prime\prime\prime})\left(\hat{Q}_{h}(x^{\prime\prime\prime},z^{\prime\prime\prime})+\mathcal{P}(\hat{m}^{\otimes 2})(x^{\prime\prime\prime},z^{\prime\prime\prime})\right)D_{qh}^{-1}(z^{\prime\prime\prime},z^{\prime\prime})G_{0}(z^{\prime\prime},z^{\prime})\right]\ ;$
@@
and take $Q_{qh}(x,z)$ for $\mathrm{i}Q_{qh}$. The accuracies are, with $\bar{s}=1$ for train and $\bar{s}=0$ for test:
@@
$\qquad\frac{1}{2}\left(1+\mathrm{erf}\left(\frac{m(1)+(\bar{s}-\rho)V_{h}(1,1)}{\sqrt{2}\sqrt{Q_{h}(1,1)-m(1)^{2}-\rho(1-\rho)V_{h}(1,1)^{2}}}\right)\right)\ .$
@@ III.2.3 Consequences
These two limits show that at finite time $t$ the GCN avoids oversmoothing and interpolates between an estimator that is only function of the features at $t=0$ and an estimator only function of the graph at $\infty. tt$ has to be fine-tuned to reach the best trade-off $t^{*}$ and the optimal performance.
@@ Acknowledgments
We acknowledge useful discussions with J. Zavatone-Veth, F. Zamponi and V. Erba. This work is supported by the Swiss National Science Foundation under grant SNSF SMArtNet (grant number 212049).
@@ Appendix A Asymptotic characterisation of the discrete GCN
The free energy of the problem is $\beta Nf=\partial_{n}\mathbb{E}_{u,\Xi,W,y}Z^{n}(n=0)$ where the partition function is
@@ A.1 Derivation of the free energy
[…]
@@
[…]
@@
II is the $n\times n$ identity and $J$ is the $n\times n$ matrix filled with ones. We introduce the $K\times K$ symmetric matrices $Q$ and $V$, filled with $(Q_{k})_{0\leq k\leq K-1}$ and $(V_{k})_{0\leq k\leq K-1}$ on the diagonal, and $(Q_{k,l})_{0\leq k<l\leq K-1}$ and $(V_{k,l})_{0\leq k<l\leq K-1}$ off the diagonal, such that $\tilde{Q}$ can be written in terms of Kronecker products as
@@
[…]
@@
$:=$ $\ \int\mathrm{d}\Theta\mathrm{d}\hat{\Theta}e^{N\phi^{(n)}(\Theta,\hat{\Theta})}\$. (160)
@@
[…]
@@
$\hat{m}\to\beta\hat{m}\$, $\hat{Q}\to\beta^{2}\hat{Q}\$, $\hat{V}\to\beta\hat{V}\$, $V\to\beta^{-1}V$ (167)
@@
Then we have the expected value of a function $g$ in $h$ $\mathbb{E}_{P_{h}}g(h)=g(h^{*})$ and the covariance $\operatorname{Cov}_{P_{h}}(h)=-\frac{1}{2}(\nabla\nabla\psi_{h}(h^{*}))^{-1}$ with $\nabla\nabla$ the Hessian; and similarly for $\mathrm{d}P_{w}$ and $\mathrm{d}P_{h}^{\prime}$.